# Tidal phenomena in the Galactic Center: The Curious Case of X7

### v1

@@ 3.2.3 Cloud structure and size
We investigated the cases of a spherical cloud uniformly sampled, a spherical cloud sampled following a three-dimensional Gaussian distribution, and an ellipsoidal cloud uniformly sampled. Here we focus on the case of the ellipsoidal cloud as it gave the most relevant results. We refer the reader to the Appendix A to see the results of the spherical cloud cases. In the ellipsoidal cloud case, we set up a cloud with particles that are uniformly distributed along a length of au, with the semi-major axis . There is negligible dependency on the other two semi-major axes ( and ), as long as they are less than . Moreover, for each particle, we set the polar angle and the azimuthal angle . These two angles have, in practice, no effect on our overall results. The initial position of the tip of the cloud is assigned from the position of S33/S0-30 in 1947. The position of the rest of the particles were assigned by adding the position of each particle from the center of the ellipsoid () to the initial position,
To assign the initial velocity of each particle, we have used a coefficient of linear increase in au yr-1 dx-1 such that there is a velocity spread from tip to tail. Moreover, we used an initial position angle of for the ridge of the ellipsoid obtained after testing angles between . In order to find the initial cloud size we tested initial lengths from 1 au to 800 au and analyzed the three dimensional separation of the tip of the simulated cloud and X7 in 2021. Figure 5 shows as a function of the initial ellipsoid length. The ellipsoidal and spherical cases are shown as dotted black and dashed red lines, respectively. In both the cases, it is possible to observe that an initial length/radius of 200 au minimizes in 2021.
In Table 1, we have summarized the different initial configurations for our simulations, their corresponding three-dimensional separation, and relative velocity in 2021. We also tested different number of particles to represent the cloud between 10 and 10000. We found that this does not affect the results in a significant way.
@@ 4 Results
In this section, we proceed to present the results of the simulations with the optimal initial conditions obtained previously. The result of one of the realizations is shown in Figure 6. This shows the sky-projected evolution of the initially ellipsoidal cloud where different colors represent the structure at different simulation times. We find that the cloud becomes more elongated through time and maintains a roughly constant position angle of 45∘ from 2002 until its pericenter passage around 2035.
4.1 X7 modeled evolution with S33/S0-30 as progenitor
Figure 7 shows the sky-projected orbits of X7 and the tip of the simulated cloud for each set of optimal velocities as dashed solid and gray dotted lines, respectively. As a reference, Sgr A* is represented as a black cross at the origin and the initial position of the tip of the cloud is displayed as a red square. The observed tip of X7 and of the simulated clouds in 2021 are shown as a purple cross and green triangles, respectively. The light blue shaded region shows the 68 confidence interval of the best-fit of the observed orbit of X7. The sky-projected orbits have a strong correspondence (by construction) with the observed orbit, the three-dimensional spatial difference of 50 au (see Table 1), is well within the uncertainty as reported by Ciurlo et al. (2023). Moreover, the best realization implies a position of the tip of the simulated cloud at 0.348 and -0.151 in 2021 in offsets from Sgr A* in right ascension and declination, respectively, both of which are within the error bars of the 2021 measurements by Ciurlo et al. (2023).
Figure 8 shows the spatial and dynamical evolution of the simulated cloud during the period of the observations from 2002 to 2021. The left-hand side panel shows the sky-projected morphology evolution of the cloud where the colors represent the particles from tip to tail. The observed orientation of X7 at different epochs is displayed as thick gray lines (the length is not to scale). As a reference, the position of Sgr A* is represented with a black cross in the upper left corner, the orbit of X7 is shown as a dashed black line, and the simulated orbit is displayed as a dashed gray line. This configuration is obtained only if the ellipsoidal cloud is initialized with a initial length 200 au, and position angle . Notice that the simulated cloud maintains a relatively constant orientation throughout the observational period from 2002 to 2021. Although the agreement with the observed orientation of X7 is not perfect, it is off by . Additionally, the length of the simulated cloud is consistent with the observed length of X7 in 2021 which is 3300 au. The right-hand side panel shows the line-of-sight velocity of the simulated cloud as a function of time. The colors represent the corresponding particles in the left panel. This analysis shows that there is a crossover of the radial velocity from tip to tail around 2018. In this case, the radial velocity of the tip decelerates by (see evolution of the black band), whereas the relative velocity of the tail remains relatively constant till 2021 (see evolution of yellow markers). It is important to remark that the observations of X7 shows exactly the same behavior although at a slightly different time (Ciurlo et al., 2023).
@@ 4.2 Comparison with observational data
Figure 9 shows a comparison of the morphological evolution of X7 observed in the L′-band and the simulated cloud during the period 2002-2021. Each panel shows an image of the sky of an area of , where Sgr A* is located in the top left corner. North and East directions are up and left, respectively. The emission of X7 is shown with contour lines and the simulated cloud is overlaid with colored markers representing the line-of-sight velocity. In 2002, X7 had a much rounder shape than the simulated cloud which is already elongated. At later times, the observations of X7 display a more elongated morphology that visually tends to align with the simulated cloud. This is clearer after 2015 when most of the observed shape of X7 is contained within the sky-projected simulated cloud. The agreement is best in the last epoch although this is not surprising since the initial velocity was selected so that tip of the cloud reproduces the position of X7 at this time.
@@
In order to make a quantitative comparison, we calculated the reduced chi-square parameter for the tip of the simulated cloud and X7 considering the three-dimensional position and velocity across every single observation epoch. Additionally, we took into account the uncertainties in the orbital fit of X7. Thus, we conducted 3000 realizations of this procedure sampling the initial conditions from the posterior distribution of the orbital fit of X7. Figure 11 shows the distribution of this procedure. This analysis also includes the impact of a hypothetical drag force due to the interaction of the cloud with the ambient medium (see Section 4.3.3). The distribution shows that the median value of and a minimum value of . Overall, the model can reproduce the observed orbit of X7 to a large extent. This result is not significantly affected by the drag force of the medium.
@@
In order to make a more appropriate comparison of the model with observational data we synthesized the Br- emission from the simulated cloud. Following Case B recombination theory, the Br- emissivity can be estimated following Schartmann et al. (2015) as
@@
The total emissivity is calculated by multiplying by , where and are the electron and proton number density of the cloud. We can obtain the corresponding flux observed from Earth by integrating over the volume element, then scaling by the inverse square law as
@@
where is the volume of the cloud, and pc is the distance to the GC. Figure 10 shows a comparison of the sky-projected observed and simulated Br- flux in 2021 as a colormap and contours, respectively. Notice that most of the simulated emission is spatially contained within the observational contours of 0.3 of the maximum emission. The simulated cloud is noticeably thinner which could be a result of not taking into account the instrument point-spread function. For completeness, a complete comparison of the evolution of the Br- flux during the period 2006-2021 is shown in Figure 18. Quantitatively, the total flux obtained from our simulated cloud is , whereas the total observed flux from X7 is in 2021 (Ciurlo et al., 2023). In summary, the model produces a cloud of similar morphology and emission of the same order of magnitude of the observations.
@@ 4.3 Secondary effects
Up to now, our models have considered solely the effect of the gravitational field of Sgr A*. In this section, we study the effects of other mechanisms that could play a role in the dynamical evolution of X7. First, we estimate the impact of stellar winds from the Wolf-Rayet (WR) stars. Within the central parsec, there are 30 WR stars with strong mass loss of 10-5 M⊙ yr-1 launched at 500-2500 km s-1 (Martins et al., 2007; Cuadra et al., 2008). Secondly, we investigate the effect of an outflow launched from the central region. The accretion flow of Sgr A* has been described successfully as a radiatively inefficient accretion flow (RIAF) due to both its inability to radiate its energy and the small fraction of the gas that ends up being accreted (e.g. Genzel et al., 2010). In this scenario, the flow gets overheated and part of it can be launched as an energetic mass outflow (Blandford & Begelman, 1999; Begelman, 2012). Assuming such an outflow indeed took place in the GC we can also estimate its impact on the evolution of X7. At last, we consider the effect of a drag force due to the interaction between the object and the ambient medium. This can alter the orbits of the test particles that represent X7 in our model. Schematic representation of these effects are shown in Figure 12.
@@ 4.3.1 Stellar winds from Wolf-Rayet stars
To date, 30% of the stars in the central parsec of the GC have complete orbital solutions (von Fellenberg et al., 2022). Thus, for the rest we used the solutions obtained by minimizing their orbital eccentricity (see Section 3.1). Then, we found the closest WR stars to X7 since the year 1947 until 2021. We considered only the five WR stars with closest approach to X7 in that period, namely E40, E48, E80, E81, and E83 (in UCLA nomenclature: S3-5, IRS13 E4, S9-9, S9-283 and S10-5, respectively; Paumard et al., 2006). These stars have winds with mass-loss rates of the order of few and velocities in the range 800-1000 (Martins et al., 2007; Cuadra et al., 2008). The ram pressure exerted by the stellar winds on X7 is given by , where is the net velocity of the wind acting on X7 taking into account the relative velocity of X7 and the WR star. The density of the stellar wind, assumed to be spherically symmetric and stationary, is denoted by and is given by,
@@
where and are the cross section in the direction of the net velocity vector and mass of X7, respectively. Then, we simulated a simple model where the net stellar wind is pointed towards the center of mass (COM) of the cloud. The results show that the motion of the simulated cloud does not change significantly. Figure 13 shows the ratio of the stellar wind and the gravitational accelerations as a function of the simulation time as a dotted-dashed blue line. Here it is possible to observe that this value remains roughly constant around . This shows that their effect is negligible during this time period.
@@ 4.3.2 Outflow from the SMBH
We performed a similar analysis as in Section 4.3.1 but considering a spherically symmetric outflow launched from the location of Sgr A*. The exact mass loss rate and velocity of the outflow were set to and cm s-1, respectively and motivated by numerical simulations of this process (e.g. Cuadra et al., 2015). Our simulation shows that the evolution of the orbit of the simulated cloud is not affected significantly by the outflow. The relative magnitude of this acceleration compared to the gravitational acceleration as a function of time is shown in Figure 13 as a dotted green line. The value is of the order of 10-17 which is even smaller than the effect of the stellar winds and, therefore, is also negligible. Moreover, even a much stronger outflow will have a negligible effect on the orbit of X7. It is relevant to mention that this result is in agreement with the analysis in Ciurlo et al. (2023), as they argued that the morphology of X7 cannot be explained by an outflow from Sgr A*.
@@ 4.3.3 The effect of a static drag force
To account for the relative motion of the ambient medium we also investigated the effect of a static drag force acting against the motion of the cloud. This has been taken into account previously for the G2 object (Madigan et al., 2017; Calderón et al., 2018), and later observed by (Gillessen et al., 2019). In principle, this would cause deviations from the Keplerian orbits of the test particles of the simulated cloud. In this case, the equation of motion becomes,
where is the gravitational constant, is the mass of Sgr A*, and are the mass and cross section of the cloud, and and are the unit vectors in the radial and velocity directions, respectively. The density of the ambient medium is considered using the model by Yuan et al. (2003) which reproduces the Chandra X-ray observations and is consistent with the latest model by Roberts et al. (2017). This density profile is given by,
where is the power-law index. Moreover, the mass of the cloud is constrained to 50 (Ciurlo et al., 2023). Then, assuming that the masses of the particles are identical, the mass of each particle is determined.
In this case, the drag force affects the orbit of the test particles. The dashed orange line in Figure 13 shows the ratio of the drag force and gravitational accelerations as a function of time. Here it can be seen that the value is 10-3 which is much larger than in the previous cases. This effect results in changes in the position of the tip of the simulated cloud that increases the three-dimensional separation of the simulated tip and X7 in 2021. Specifically, the fiducial model results in a separation of 50 au while the calculations with the drag force shows it to be 130 au. However, there is no significant change in the length of the cloud. Although increases, the result is still within the margin of the errors. Thus, this effect does not impact significantly the overall results. It is to be noted that a non-static drag force could have slightly different effects (Madigan et al., 2017).
@@ 5 Discussion
We have shown that ejecta from S33/S0-30 provided with an ejection velocity of 610 in 1947 can reproduce the orbit of X7 such that the tip of the simulated cloud and the observed tip of X7 in 2021 have a separation, , of around 50 au. We have also shown that an initially elongated cloud with an initial velocity gradient can reproduce the observed orientation of X7 (see Figure 8). In this section, we interpret these results based on our proposed hypotheses for the origin of X7.
@@ 5.1 A star in an extensive mass-losing phase
We started with the hypothesis that X7 formed when a star in the past 200 years went through a phase of excessive mass loss such as a Luminous-Blue-Variable phase. The abundance of O-type and WR stars in the galactic center support this idea, since it is possible for a O-type star to go through a LBV phase before becoming a WR star (e.g. Crowther, 2007). The WR stars in the GC have a mass loss rate of a few (Martins et al., 2007). With this rate, it would take approximately 15 years to lose an amount of mass comparable to the mass of X7. However, the best candidate to be the origin of X7, S33/S0-30, is a B-type star whose mass-loss rates are not as high (10; Martins et al., 2008). Even if the mass-loss rates were to go an order or two higher, the timescale needed is very long to generate a single coherent structure comparable to X7. Moreover, such a mass-loss will occur in a spherically-symmetric manner, whereas our model favors an ellipsoidal ejecta. Based on this, it is very unlikely for X7 to have originated in this way.

### v2

@@ 3.2.3 Cloud structure and size
We investigated the cases of a spherical cloud uniformly sampled, a spherical cloud sampled following a three-dimensional Gaussian distribution, and an ellipsoidal cloud uniformly sampled. Here we focus on the case of the ellipsoidal cloud as it gave the most relevant results. We refer the reader to Appendix A to see the results of the spherical cloud cases. In the ellipsoidal cloud case, we set up a cloud with particles that are uniformly distributed along a length of au, with the semimajor axis . There is negligible dependence on the other two semimajor axes ( and ), as long as they are less than . Moreover, for each particle, we set the polar angle and the azimuthal angle . These two angles have, in practice, no effect on our overall results. The initial position of the tip of the cloud is assigned from the position of S33/S0-30 in 1947. The position of the rest of the particles were assigned by adding the position of each particle from the center of the ellipsoid () to the initial position,
To assign the initial velocity of each particle, we used a coefficient of linear increase () in au yr-1 dx-1 such that there is a velocity spread of 100 from tip to tail. Moreover, we used an initial position angle of for the ridge of the ellipsoid obtained after testing angles in the range . To find the initial cloud size, we tested initial lengths from 1 au to 800 au and analyzed the three dimensional separation of the tip of the simulated cloud and X7 in 2021. Figure 5 shows as a function of the initial ellipsoid length for the ellipsoidal and spherical cases. In both the cases, it is possible to observe that an initial length/radius of 200 au minimizes in 2021.
In Table 1, we have summarized the different initial configurations for our simulations, their corresponding three-dimensional separation, and relative velocity in 2021. We also tested different number of particles to represent the cloud (), between 10 and 10000. We find that this does not affect the results in a significant way.
@@ 4 Results
In this section we present the results of the simulations with the optimal initial conditions obtained previously. The result of one of the realizations is presented in Fig. 6, which shows the sky-projected evolution of the initially ellipsoidal cloud at different simulation times. We find that the cloud becomes more elongated through time and maintains a roughly constant position angle of 45∘ from 2002 until its pericenter passage around 2035.
4.1 X7 modeled evolution with S33/S0-30 as the progenitor
Figure 7 shows the sky-projected orbits of X7 and the tip of the simulated cloud for each set of optimal velocities. The sky-projected orbits have a strong correspondence (by construction) with the observed orbit, the three-dimensional spatial difference of 50 au (see Table 1), is well within the uncertainty as reported by Ciurlo et al. (2023). Moreover, the best realization implies a position of the tip of the simulated cloud at 0.348 and -0.151 in 2021 in offsets from Sgr A* in right ascension and declination, respectively, both of which are within the error bars of the 2021 measurements by Ciurlo et al. (2023).
Figure 8 shows the spatial and dynamical evolution of the simulated cloud during the period of the observations from 2002 to 2021. The left panel shows the sky-projected morphology evolution of the cloud. This configuration is obtained only if the ellipsoidal cloud is initialized with a initial length 200 au, and position angle . Notice that the simulated cloud maintains a relatively constant orientation throughout the observational period from 2002 to 2021. Although the agreement with the observed orientation of X7 is not perfect, it is off by . Additionally, the length of the simulated cloud is consistent with the observed length of X7 in 2021, which is 3300 au. The right panel shows the line-of-sight velocity of the simulated cloud as a function of time. This analysis shows that there is a crossover of the radial velocity from tip to tail around 2018. In this case, the radial velocity of the tip decelerates by (see the evolution of the black band), whereas the relative velocity of the tail remains relatively constant till 2021 (see the evolution of the yellow markers). It is important to remark that the observations of X7 shows exactly the same behavior although at a slightly different time (Ciurlo et al., 2023).
@@ 4.2 Comparison with observational data
Figure 9 shows a comparison of the morphological evolution of X7 observed in the L′ band and the simulated cloud during the period 2002-2021. Each panel shows an image of the sky of an area of , where Sgr A* is located in the top left corner. In 2002, X7 had a much rounder shape than the simulated cloud, which is already elongated. At later times, the observations of X7 display a more elongated morphology that visually tends to align with the simulated cloud. This is clearer after 2015 when most of the observed shape of X7 is contained within the sky-projected simulated cloud. The agreement is best in the last epoch although this is not surprising since the initial velocity was selected so that tip of the cloud reproduces the position of X7 at this time.
@@
To make a quantitative comparison, we calculated the reduced chi-square parameter () for the tip of the simulated cloud and X7 considering the three-dimensional position and velocity across every single observation epoch. Additionally, we took into account the uncertainties in the orbital fit of X7. Thus, we conducted 3000 realizations of this procedure sampling the initial conditions from the posterior distribution of the orbital fit of X7. Figure 10 shows the distribution of this procedure. This analysis also includes the impact of a hypothetical drag force due to the interaction of the cloud with the ambient medium (see Sect. 4.3.3). The distribution shows that the median value of and a minimum value of . Overall, the model can reproduce the observed orbit of X7 to a large extent. This result is not significantly affected by the drag force of the medium.
@@
To make a more appropriate comparison of the model with observational data, we synthesized the Br emission from the simulated cloud. Following Case B recombination theory, the Br emissivity can be estimated following Schartmann et al. (2015) as
@@
We assumed a temperature of K for the simulated cloud, which is typical for gas structures in the central parsec due to the strong ultraviolet radiation from the young, massive stars (Zhao et al., 2009; Gillessen et al., 2012). The total emissivity was calculated by multiplying by , where and are the electron and proton number density of the cloud. We obtained the corresponding flux observed from Earth by integrating over the volume element and then scaling it by the inverse square law as
@@
where is the volume of the cloud, and pc is the distance to the GC. Figure 11 shows a comparison of the sky-projected observed and simulated Br flux as seen from Earth in 2021. Notice that most of the simulated emission is spatially contained within the observational contours of 0.3 of the maximum emission. The simulated cloud is noticeably thinner, which could be a result of not taking into account the instrument point-spread function. For completeness, a complete comparison of the evolution of the Br flux during the period 2006-2021 is shown in Fig. 18. Quantitatively, the total flux obtained from our simulated cloud is , whereas the total observed flux from X7 is in 2021 (Ciurlo et al., 2023). In summary, the model produces a cloud of similar morphology and emission of the same order of magnitude of the observations.
@@ 4.3 Secondary effects
Up to now, our models have considered solely the effect of the gravitational field of Sgr A*. In this section we study the effects of other mechanisms that could play a role in the dynamical evolution of X7. First, we estimated the impact of stellar winds from the WR stars. Within the central parsec, there are 30 WR stars with strong mass loss of 10-5 M⊙ yr-1 launched at 500-2500 km s-1 (Martins et al., 2007; Cuadra et al., 2008). Secondly, we investigated the effect of an outflow launched from the central region. The accretion flow of Sgr A* has been described successfully as a radiatively inefficient accretion flow due to both its inability to radiate its energy and the small fraction of the gas that ends up being accreted (e.g., Genzel et al., 2010). In this scenario, the flow gets overheated and part of it can be launched as an energetic mass outflow (Blandford & Begelman, 1999; Begelman, 2012). Assuming such an outflow indeed took place in the GC, we can also estimate its impact on the evolution of X7. Lastly, we considered the effect of a drag force due to the interaction between the object and the ambient medium. This can alter the orbits of the test particles that represent X7 in our model. Schematic representation of these effects are shown in Fig. 12.
@@ 4.3.1 Stellar winds from Wolf-Rayet stars
To date, 18% of the stars in the central parsec of the GC have complete orbital solutions (von Fellenberg et al., 2022). For the rest, we used the solutions obtained by minimizing their orbital eccentricity (see Sect. 3.1). Then we found the closest WR stars to X7 since the year 1947 until 2021. We considered only the five WR stars with closest approach to X7 in that period, namely E40, E48, E80, E81, and E83 (in UCLA nomenclature: S3-5, IRS13 E4, S9-9, S9-283, and S10-5, respectively; Paumard et al., 2006). These stars have winds with mass-loss rates on the order of a few times and velocities in the range 800-1000 (Martins et al., 2007; Cuadra et al., 2008). The ram pressure exerted by the stellar winds on X7 is given by , where is the net velocity of the wind acting on X7 taking into account the relative velocity of X7 and the WR star. The density of the stellar wind, assumed to be spherically symmetric and stationary, is denoted by and is given by
@@
where and are the cross section in the direction of the net velocity vector and mass of X7, respectively. Then, we simulated a simple model where the net stellar wind is pointed toward the center of mass of the cloud. The results show that the motion of the simulated cloud does not change significantly. Figure 13 shows the ratio of the stellar wind and the gravitational accelerations as a function of the simulation time. Here it is possible to observe that this value remains roughly constant around . This shows that their effect is negligible during this time period.
@@ 4.3.2 Outflow from the SMBH
We performed a similar analysis as in Sect. 4.3.1 but considering a spherically symmetric outflow launched from the location of Sgr A*. The exact mass-loss rate and velocity of the outflow were set to and cm s-1, respectively and motivated by numerical simulations of this process (e.g., Cuadra et al., 2015). Our simulation shows that the evolution of the orbit of the simulated cloud is not affected significantly by the outflow. The relative magnitude of this acceleration compared to the gravitational acceleration as a function of time is shown in Fig. 13. The value is on the order of 10-17, which is even smaller than the effect of the stellar winds and, therefore, is also negligible. Moreover, even a much stronger outflow will have a negligible effect on the orbit of X7. It is relevant to mention that this result is in agreement with the analysis in Ciurlo et al. (2023), as they argued that the morphology of X7 cannot be explained by an outflow from Sgr A*.
@@ 4.3.3 The effect of a static drag force
To account for the relative motion of the ambient medium we also investigated the effect of a static drag force acting against the motion of the cloud. This has been taken into account previously for the G2 object (Madigan et al., 2017; Calderón et al., 2018), and later observed by (Gillessen et al., 2019). In principle, this would cause deviations from the Keplerian orbits of the test particles of the simulated cloud. In this case, the equation of motion becomes
where is the gravitational constant, is the mass of Sgr A*, and are the mass and cross section of the cloud, and and are the unit vectors in the radial and velocity directions, respectively. The density of the ambient medium () is based on the model from Yuan et al. (2003), which reproduces the Chandra X-ray observations and is consistent with the latest model by Roberts et al. (2017). This density profile is given by
where is the power-law index set to based on the analysis of Burkert et al. (2012), although we also tested and and found no substantive change in the result. Moreover, the mass of the cloud is constrained to 50 (Ciurlo et al., 2023). Then, assuming that the masses of the particles are identical, the mass of each particle was determined.
In this case, the drag force affects the orbit of the test particles. Figure 13 shows the ratio of the drag force and gravitational accelerations as a function of time. Here it can be seen that the value is 10-3, which is much larger than in the previous cases. This effect results in changes in the position of the tip of the simulated cloud that increases the three-dimensional separation of the simulated tip and X7 in 2021. Specifically, the fiducial model results in a separation of 50 au while the calculations with the drag force shows it to be 130 au. However, there is no significant change in the length of the cloud. Although increases, the result is still within the margin of the errors. Thus, this effect does not impact significantly the overall results. It is to be noted that a non-static drag force could have slightly different effects (Madigan et al., 2017).
@@ 5 Discussion
We have shown that ejecta from S33/S0-30 provided with an ejection velocity of 610 in 1947 can reproduce the orbit of X7 such that the tip of the simulated cloud and the observed tip of X7 in 2021 have a separation () of around 50 au. We have also shown that an initially elongated cloud with an initial velocity gradient can reproduce the observed orientation of X7 (see Fig. 8). In this section we interpret these results based on our proposed hypotheses for the origin of X7.
@@ 5.1 A star in an extensive mass-losing phase
We started with the hypothesis that X7 formed when a star in the past 200 years went through a phase of excessive mass loss such as a LBV phase. The abundance of O-type and WR stars in the GC support this idea, since it is possible for a O-type star to go through a LBV phase before becoming a WR star (e.g., Crowther, 2007). The WR stars in the GC have a mass-loss rate of a few times (Martins et al., 2007). With this rate, it would take approximately 15 years to lose an amount of mass comparable to the mass of X7. However, the best candidate to be the origin of X7, S33/S0-30, is a B-type star whose mass-loss rates are not as high (10; Martins et al., 2008). Even if the mass-loss rates were to go an order or two higher, the timescale needed would be too long to generate a single coherent structure comparable to X7. Moreover, such a mass loss will occur in a spherically symmetric manner, whereas our model favors an ellipsoidal ejecta. Based on this, it is very unlikely that X7 originated in this way.